# Data-efficient Targeted Token-level Preference Optimization for LLM-based Text-to-Speech

### v1

@@ Abstract
Aligning text-to-speech (TTS) system outputs with human feedback through preference optimization has been shown to effectively improve the robustness and naturalness of language model-based TTS models. Current approaches primarily require paired desirable and undesirable samples at the utterance level. However, such pairs are often limited in TTS output data, and utterance-level formulation prevents fine-grained token-level optimization needed for accurate pronunciation alignment. In this study, we propose TKTO that eliminates the need for paired data, enabling a more data-efficient training paradigm, and directly targets token-level units, automatically providing fine-grained alignment signals without token-level annotations. TKTO improves the challenging Japanese TTS accuracy by 39% and reduces CER by 54%, automatically assigning $12.8\times$ stronger reward to targeted tokens.
@@ 1 Introduction
Recent advances in neural network technology have enabled high-fidelity text-to-speech (TTS) synthesis models Chen et al. (2025); Meng et al. (2025); Li et al. (2024). They typically convert input text into a phoneme sequence using a grapheme-to-phoneme (G2P) converter Oura et al. (2010), followed by generative models from the phoneme sequence Wang et al. (2025b); Nishimura et al. . However, G2P is generally based on morphological analysis, and thus may fail to generate correct pronunciations in ambiguous languages like Japanese, where the reading and meaning of words can change depending on context (Figure 1).
@@
  • Problem formulation: we pioneer the challenging task of ambiguous Japanese pronunciation as a token-level preference optimization problem.
@@
  • Wide evaluation: we demonstrate that TKTO not only reduces character error rate (CER) by 54% but also improves the accuracy of Japanese pronunciation by 39 %, surpassing industry models. It selectively increases the generation probability of desirable tokens and assigns 12.8× stronger rewards to targeted tokens.
@@ 3.1 Targeted Token Weight Estimation
We propose a KTO-based approach to construct contrastive LLMs. KTO Ethayarajh et al. (2024) uses unpaired data, enabling more flexible training. We build two contrastive LLMs, $\pi^{+}$ and $\pi^{-}$, where $\pi^{+}$ favors high- and $\pi^{-}$ favors low-reward tokens. We train $\pi^{+}$ with normal labels and obtain $\pi^{-}$ by simply flipping the labels, treating desirable as undesirable and vice versa.
@@
We estimate the token’s weight Liu et al. (2025) as:
@@
where $\log\frac{\pi^{+}(y_{t}\mid x,y_{<t})}{\pi^{-}(y_{t}\mid x,y_{<t})}$ estimates the token’s reward (Rafailov et al., 2024). We clamp this reward between $L$ and $U$ to enhance optimization stability. For desirable samples, we set $\mu>0$; for undesirable ones, we set $\mu<0$.
@@ 3.2 Token-level KTO
We define the reward for each token $y_{t}$ as its log-ratio against a reference policy $\pi_{\text{ref}}$:
@@
where $\sigma(\cdot)$ is the sigmoid function, $\beta$ controls curvature, and $\lambda_{D}$, $\lambda_{U}$ adjust aversion for desirable and undesirable samples.
@@
The total TKTO loss is defined by summing token-level values across the sequence and weighting them by an importance weight $w_{t}$:
@@ 4.1 Experimental Setup
We use CER to evaluate overall robustness and the bad case ratio (Bad), where samples with CER exceed 0.3.
@@
We employ the naturalness mean opinion score (NMOS) to evaluate the naturalness. In addition, we conduct an ABX test. Please see Appendix A for more details.
@@ 4.2 Objective Evaluation Results
Main Results.
Table 1 presents the objective evaluation results. Our TKTO model achieves the highest Japanese TTS accuracy and the lowest CER and bad case ratio. The non-LLM baseline F5-TTS shows extremely low accuracy regardless of whether it uses G2P or not, highlighting the importance of LLMs to consider contextual information. The use of unpaired data allows us to leverage samples that are always desirable or always undesirable, which cannot be used in paired training. This leads to a significant improvement in accuracy (0.949 or 0.958), surpassing even strong industry models. Using only paired data excludes samples that always produce desirable outputs, causing the model to overfit toward undesirable examples. As a result, accuracy improvement is limited, although the CER and bad case ratio become slightly lower. Compared with KTO, our model achieves improvements across all metrics, demonstrating the effectiveness of token-level targeted optimization.
Training Dynamics.
Token Weight Analysis.
Figure 4 illustrates the token rewards $\log\frac{\pi^{+}(y_{t}\mid x,y^{<t})}{\pi^{-}(y_{t}\mid x,y^{<t})}$ for desirable and undesirable tokens for the target character "辛" and the overall average. The desirable tokens have a higher reward (0.22) compared to the overall mean (0.12), while the undesirable tokens show a much lower value (-1.54), indicating that targeted tokens are automatically assigned larger weights. For undesirable samples, two peaks are observed when a clear difference between $\pi^{+}$ and $\pi^{-}$ emerges; the reward drops sharply, whereas when no significant difference exists. Figure 5 also presents a case study showing that the tokens corresponding to the target characters receive notably higher weights.
@@ 5 Conclusion
The evaluations are conducted on only challenging Japanese data. Although multiple preference optimization methods and configurations were tested, the base model was limited to CosyVoice2 0.5B model due to the computational resources. While this work only focuses on preference optimization for TTS, the proposed TKTO framework can potentially be applied to other text generation tasks where specific tokens play a crucial role.
@@
Appendix A Subjective Evaluation Details
The instructions for the subjective evaluation are as follows. A Japanese native human annotator listened to the speech samples and rated their naturalness as NMOS on a 5-point scale, where 1 indicates very unnatural and 5 indicates completely natural. In addition, we conducted an ABX test in which the participant listened to two generated speech samples from different models but based on the same input and then chose the one that sounds more natural; if the samples are too similar to distinguish, they are instructed to indicate a tie. The annotator was recruited via a crowdsourcing platform and received appropriate compensation for their work.
@@ Appendix B Implementation Details
Appendix C Parameter Sensitivity Analysis

### v2

@@ Abstract
Aligning text-to-speech (TTS) system outputs with human feedback through preference optimization has been shown to effectively improve the robustness and naturalness of LLM-based TTS models. Current approaches primarily require paired desirable and undesirable samples at the utterance level. However, such pairs are often limited in TTS output data, and utterance-level formulation prevents fine-grained token-level optimization needed for accurate pronunciation alignment. In this study, we propose TKTO that eliminates the need for paired data, enabling a more data-efficient training paradigm, and directly targets token-level units, automatically providing fine-grained alignment signals without token-level annotations. TKTO improves the challenging Japanese TTS accuracy by 39% and reduces CER by 54%, leveraging $6\times$ more training data and assigning $12.8\times$ stronger reward to targeted tokens.
@@ 1 Introduction
Recent advances in neural network technology have enabled high-fidelity text-to-speech (TTS) synthesis models Chen et al. (2025); Meng et al. (2025); Li et al. (2024). They typically convert input text into a phoneme sequence using a grapheme-to-phoneme (G2P) converter Oura et al. (2010), followed by generative models from the phoneme sequence Wang et al. (2025b); Nishimura et al. (2025). However, G2P is generally based on morphological analysis, and thus may fail to generate correct pronunciations in ambiguous languages like Japanese, where the reading and meaning of words can change depending on context (Figure 1).
@@
  • Problem formulation: we pioneer the challenging task of ambiguous pronunciation as a token-level preference optimization problem.
@@
  • Wide evaluation: we demonstrate that TKTO reduces character error rate (CER) by 54% and improves the accuracy of Japanese pronunciation by 39 %. By leveraging unpaired data, TKTO can utilize $6\times$ more training data than DPO, leading to higher data efficiency. It also automatically assigns 12.8× stronger rewards to targeted tokens.
@@ 3.1 Targeted Token Weight Estimation
We propose a KTO-based approach to construct contrastive LLMs. KTO Ethayarajh et al. (2024) uses unpaired data, enabling more flexible training. Details are provided in Appendix A. We build two contrastive LLMs, $\pi^{+}$ and $\pi^{-}$, which do not share parameters. We train $\pi^{+}$ using the original desirable/undesirable labels, and obtain $\pi^{-}$ by training on the same data with the labels swapped, i.e., treating desirable samples as undesirable and vice versa.
@@
We estimate the token’s weight $w_{t}$ Liu et al. (2025) as:
@@
where $\log\frac{\pi^{+}(y_{t}\mid x,y_{<t})}{\pi^{-}(y_{t}\mid x,y_{<t})}$ estimates the token’s reward (Rafailov et al., 2024). $L<U$ are lower and upper clamping bounds to improve optimization stability. $\mu$ is a scaling factor that controls the strength of reward weighting. For desirable samples, we set $\mu>0$; for undesirable ones, we set $\mu<0$.
@@ 3.2 Token-level KTO
We define the reward $r_{\theta,t}(x,y)$ for each token $y_{t}$ as its log-ratio against a reference policy $\pi_{\text{ref}}$:
@@
where $\sigma(\cdot)$ is the sigmoid function, $\beta$ controls curvature, and $\lambda_{D}$, $\lambda_{U}$ adjust aversion weights.
@@
The TKTO loss is a weighted sum of token-level values, with importance weights $w_{t}$:
@@ 4.1 Experimental Setup
Accuracy measures the proportion of samples in which the generated pronunciation matches the correct reading for a target ambiguous word (e.g., "辛い" ). We also use CER to evaluate overall robustness and the bad case ratio (Bad), where samples with CER exceed 0.3.
@@
We employ the naturalness mean opinion score (NMOS) to evaluate the naturalness. point scale, where 1 indicates very unnatural and 5 indicates completely natural. In addition, we conduct an ABX test. The instructions for the subjective evaluation are as follows. Three Japanese native human annotators listened to the speech samples and rated their naturalness as NMOS on a 5-point scale, where 1 indicates very unnatural and 5 indicates completely natural. In addition, we conducted an ABX test in which the participant listened to two generated speech samples from different models but based on the same input and then chose the one that sounds more natural; if the samples are too similar to distinguish, they are instructed to indicate a tie. The annotator was recruited via a crowdsourcing platform and received appropriate compensation for their work.
@@ 4.2 Objective Evaluation Results
(a) TKTO (Ours)
(b) SFT
Main Results on Japanese.
Table 1 presents the objective evaluation results. Our TKTO model achieves the highest Japanese TTS accuracy and the lowest CER and bad case ratio. The non-LLM baseline F5-TTS shows extremely low accuracy regardless of whether it uses G2P or not, highlighting the importance of LLMs to consider contextual information. By leveraging unpaired data, we can utilize up to $6\times$ more training samples than paired methods. This leads to a significant improvement in accuracy (0.668→0.958), surpassing even strong industry models. Paired training relies on a very narrow and biased subset that represents only rare borderline cases where both correct and incorrect pronunciations appear for the same text. This limits the learning signal and reduces generalization. Compared with KTO, our model achieves improvements across all metrics, demonstrating the effectiveness of token-level targeted optimization.
Training Dynamics.
Token Weight Analysis.
Figure 4 illustrates the token rewards $\log\frac{\pi^{+}(y_{t}\mid x,y^{<t})}{\pi^{-}(y_{t}\mid x,y^{<t})}$ for desirable and undesirable tokens for the target character "辛" and the overall average. The desirable tokens have a higher reward (0.22) compared to the overall mean (0.12), while the undesirable tokens show a much lower value (-1.54), indicating that targeted tokens are automatically assigned larger weights. For undesirable samples, two peaks are observed when a clear difference between $\pi^{+}$ and $\pi^{-}$ emerges; the reward drops sharply, whereas when no significant difference exists.
Case Study.
Figure 5 presents a case study showing that the tokens corresponding to the target characters receive notably higher weights compared to non-targeted tokens.
@@ 4.3 Subjective Evaluation Results
4.4 Computational Cost and Data Efficiency
Building Two Contrastive LLMs.
Our TKTO framework requires training two additional contrastive LLMs to estimate token weights, which introduces an extra pre-computation step in the pipeline. However, the additional training cost is relatively small: in our setup, training these two contrastive LLMs takes approximately 10 minutes on 8×A100 GPUs. This overhead is negligible compared to the overall TTS training process, where pre-training on tens of thousands of hours of data typically requires weeks of computation.
Data Efficiency with Limited Pairs.
In our dataset, only 10.5% of utterances contain both a desirable and an undesirable output, while 89.5% contain only one side. This imbalance occurs because modern TTS systems tend to produce consistently correct or incorrect pronunciations, making paired samples rare. As a result, paired methods such as DPO can use only 1.5K samples, whereas unpaired methods can utilize 9K samples. In our experiments, TKTO is able to leverage the unpaired samples, resulting in approximately 6× more usable data than what is accessible to paired preference optimization methods.
4.5 Additional Validation on Chinese
In addition to Japanese, we evaluate the generalization capability of our method on Chinese. Following the same experimental setup as in the Japanese evaluation, we generate 5,000 sentences containing the polyphonic character "行" which has two context-dependent readings, “xíng” and “háng.” For each sentence, we generate five speech samples using the base TTS model with a Chinese speaker, resulting in 1.5K paired samples and 4K unpaired samples. We evaluate the correctness of each reading using the wav2vec2-mms-1b-cmn-phonetic checkpoint [^1] https://huggingface.co/Chuatury/wav2vec2-mms-1b-cmn-phonetic.
Results.
Table 3 presents the Chinese evaluation results. TKTO achieves the highest accuracy and the lowest CER and bad-case ratio among all compared methods. These results demonstrate that TKTO generalizes beyond Japanese and is effective for polyphonic disambiguation in Chinese.
Moethods PO Data Acc $\uparrow$ CER $\downarrow$ Bad $\downarrow$
Base model 0.645 0.024 0.010
DPO Paired 0.682 0.021 0.007
KTO Paired 0.712 0.026 0.009
KTO Unpaired 0.832 0.018 0.010
TKTO (ours) Paired 0.746 0.024 0.007
TKTO (ours) Unpaired 0.898 0.014 0.003
Table 3: Results on Chinese polyphonic disambiguation.
@@ 5 Conclusion
The evaluations are conducted on only challenging Japanese and Chinese data. Although multiple preference optimization methods and configurations were tested, we only evaluate a single model size (0.5B), as larger versions of the same backbone are not publicly available and computational constraints prevent extensive pretraining and scaling experiments. Investigating scaling behavior and extending the approach to other model architectures remain to be explored. While this work only focuses on preference optimization for TTS, the proposed TKTO framework can potentially be applied to other text generation tasks where specific tokens play a crucial role.
Acknowledgments
The authors would like to sincerely thank the anonymous reviewers for their valuable comments and helpful suggestions.
@@ Acknowledgments
Appendix A Kahneman–Tversky Optimization
Kahneman–Tversky Optimization (KTO) Ethayarajh et al. (2024) is a pair-free alignment objective that directly maximizes a proxy of human utility using binary feedback indicating whether an output is desirable or undesirable. Unlike DPO Rafailov et al. (2023), KTO does not require paired comparisons.
Reference-adjusted reward.
Given an input $x$, output $y$, current policy $\pi_{\theta}$, and reference policy $\pi_{\text{ref}}$, we define the reward $r_{\theta}(x,y)$ as:
$r_{\theta}(x,y)=\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}.$ (7)
A reference point $z_{0}$ is defined as:
$z_{0}=\mathrm{KL}\!\left(\pi_{\theta}(\cdot\mid x)\,\|\,\pi_{\text{ref}}(\cdot\mid x)\right),$ (8)
which captures the expected reward.
Value function.
To model loss aversion and diminishing sensitivity, KTO uses a logistic value function:
$v(x,y)=\begin{cases}\lambda_{D}\,\sigma\!\left(\beta(r_{\theta}(x,y)-z_{0})\right),\\ \quad\quad\quad\text{if }y\sim y_{\text{desirable}}\mid x,\\ \lambda_{U}\,\sigma\!\left(\beta(z_{0}-r_{\theta}(x,y))\right),\\ \quad\quad\quad\text{if }y\sim y_{\text{undesirable}}\mid x,\end{cases}$ (9)
where $\sigma(\cdot)$ is the sigmoid function, $\beta$ controls risk sensitivity, and $\lambda_{D},\lambda_{U}$ control gains and losses.
KTO objective.
The KTO loss is defined as:
$\mathcal{L}_{\text{KTO}}(\pi_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}}\big[\,-v(x,y)\,\big],$ (10)
Connection to TKTO.
Our TKTO extends KTO by lifting the optimization from the sequence level to the token level for TTS. While KTO assigns a single utility to an entire output $y$, TKTO introduces token-wise weights that estimate the relative contribution of each token to human utility. By explicitly modeling heterogeneous token importance, TKTO enables fine-grained credit assignment across tokens, allowing the model to emphasize critical reasoning or decision-making steps while down-weighting less informative tokens. In this sense, TKTO can be viewed as a token-wise generalization of KTO that preserves its prospect-theoretic inductive bias while substantially improving optimization granularity.
@@ Appendix C Parameter Sensitivity Analysis
Female Male
Model CER $\downarrow$ Bad $\downarrow$ CER $\downarrow$ Bad $\downarrow$
Base model 0.152 0.110 0.151 0.094
KTO 0.072 0.050 0.076 0.047
TKTO (ours) 0.065 0.044 0.069 0.045
Table 5: Evaluation results using the nvidia/parakeet-tdt_ctc-0.6b-ja ASR model.
@@
Appendix D Robustness across ASR Backends
To verify the robustness of our evaluation, we conduct additional experiments using the parakeet-tdt_ctc-0.6b-ja [^2] https://huggingface.co/nvidia/parakeet-tdt_ctc-0.6b-ja ASR model with paired preference data. Table 5 shows consistent improvements over the baselines, mirroring the trend observed with Whisper. As shown in Table 5, TKTO achieves the lowest CER and Bad ratios for both speakers, confirming that the performance gains are not tied to a particular ASR backend.